# Cohomology of Lie algebroids over Algebraic spaces

### v7

@@ 3.3. Chern classes:
Lie algebroid (or Lie-Rinehart algebra) connections, curvatures and the associated Chern classes (one of the Characteristic classes) has been studied in the context of differential geometry [10] and algebric geometry [23, 24, 25]. We review and extend some of the facts in our context as follows.
@@
where $Ext^{1}(L,E,\nabla)$ is the set of equivalence classes of extensions of $L$ by the flat connection $(E,\nabla)$ [25]. This result can be viewed in the algebraic geometry set up (over the affine scheme $SpecR$). Analogously, for a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$ we get the isomorphism of $\mathbb{K}$-vector spaces
@@
for any two sections $D,D^{\prime}$ of $\mathcal{T}_{X}$. For a $\mathcal{L}$-connection $\nabla$ on $\mathcal{E}$, one verifies that the curvature $R_{\nabla}\in\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{2}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E}))$, where $\mathcal{L}$ is a Lie algebroid.
@@
$c_{1}(\mathcal{E})=Trace(R_{\nabla})\in\mathbb{H}^{2}(X,\mathcal{O}_{X})$
where $Trace:\mathcal{E}nd(\mathcal{E})\rightarrow\mathcal{O}_{X}$ is the trace map.
The restriction of a $\mathcal{T}_{X}$-connection to the Lie algebroid $\mathcal{L}:=\mathcal{T}_{X}(-logY)$ ((((i.e. $\Omega_{\mathcal{L}}^{1}:=\Omega_{X}^{1}(logY)$) induces a logarithmic connection which is a meromorphic connection with poles along the divisor $Y$. Moreover, if $Y$ is a free divisor, then for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank with a $\mathcal{L}$-connection we get the first Chern class $c_{1}(\mathcal{E})\in\mathbb{H}^{2}(\mathcal{L},\mathcal{O}_{X})$. Note that a $\mathcal{L}$-connection might not exist in general. We can extend the notion of Chern classes for $\mathcal{T}_{X}$ or $\mathcal{T}_{X}(-logY)$ to arbitary Lie algebroids.
@@ 4. Logarithmic de Rham cohomology
We compute algebraic (analytic) de Rham cohomology groups of a family of non-singular varieties $Y_{t}$ for $t\in\mathbb{C}\setminus\{0\}$ and their associated singular variety $Y:=Y_{0}$ (which appears as principal divisors [29]), i.e. cohomology of the Lie algebroid $\mathcal{T}_{Y_{t}}$ for $t\in\mathbb{C}$. Then, we compute hypercohomology of the logarithmic de Rham complexes [27], which simplifies to the Lie algebroid cohomology for $\mathcal{T}_{X}(-logY)$ [32], as we work over some free divisors $Y\subset X$ [4, 32].
@@
(using the standard Lemma of Atiyah-Hodge [36]), similar results appear in [4].
@@ 5. Hochschild Cohomology for a Lie algebroid over $a$-spaces
In this section, we define Hochschild cohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra (see Section 2.3), using (Hopf-)Hochschild cohomology of a left bialgebroid [19] as local descriptions. In particular, we describe the cases associated with universal enveloping algebroid and jet algebroid of a locally free Lie algebroid over an $a$-space. The cohomology groups can be computed using suitable standard complexes. We prove a version of Hochschild-Kostant-Rosenberg (HKR) theorem and dual HKR theorem. It is done by following the local counterpart as can be found in [19], where the authors proved (dual) HKR theorem for (finitely generated) projective Lie-Rinehart algebras. Moreover, we present the HKR theorem for the sheaf of logarithmic differential operators and the sheaf of noncommuative differential operators.
@@ 5.1. Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra
where $C^{\bullet}(\mathcal{A}(U))=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}},b)$ is the Hochschild cochain complex of the $\mathcal{O}_{X}(U)/{\mathbb{K}}$-bialgebra $\mathcal{A}(U)$ with the differential $b$ defined for any $a_{1},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
$b(a_{1}\otimes\cdots\otimes a_{n})=1\otimes a_{1}\otimes\cdots\otimes a_{n}+\sum^{n}_{i=1}(-1)^{i}a_{1}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}$ $+(-1)^{n+1}a_{1}\otimes\cdots\otimes a_{n}\otimes 1.$
@@
where $Cob^{\bullet}({\mathcal{A}(U)}):=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}},b^{\prime})$ is the cobar complex for $\mathcal{O}_{X}(U)$ in the category of left $\mathcal{A}(U)$-comodules with the differential $b^{\prime}$ defined for any $a_{0},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
@@
Now, we describe Hochschild cohomology for some special $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras, namely for the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [29] and the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ [29] of a (locally free) Lie algebroid $\mathcal{L}$ over a special $a$-space (as mentioned in Note 2.2) $(X,\mathcal{O}_{X})$.
@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
The sheaf of $\mathcal{L}$-poly-vector fields on $X$ is defined as $\mathcal{T}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\wedge^{i}\mathcal{L}$. This generalizes the space of multisections of the tangent bundle (i.e. multi-vector fields) descried for different $a$-spaces occurs in geometries. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. For the case of a smooth manifold $X$, Kontsevich introduced the (sheaf of) poly-differential operators on $X$. It induces a subcomplex of the Hochschild cochain complex for $\mathcal{O}_{X}$. Its analogue in the context of Lie algebroid is the sheaf of $\mathcal{L}$-poly differential operators $\mathcal{D}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes_{\mathcal{O}_{X}}^{i}}$. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. In particular, when $\mathcal{L}=\mathcal{T}_{X}$ we denote these sheaves by the standard notions $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ respectively. The so-called HKR map is a quasi-isomorphism between $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ (of differential graded $\mathbb{K}_{X}$-vector spaces). See [7] for more generalities.
@@
We know that the sheaf of symmetric algebras $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ can be viewed as universal enveloping algebroid of the Lie algebroid $\mathcal{L}$ if the $\mathcal{O}_{X}$-module $\mathcal{L}$ is equipped with zero bracket and zero anchor (on each space of sections) [29]. Recall that we have PBW presheaf homomorphism of presheaf of $\mathcal{O}_{X}$-algebas given as
$\theta:S_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$
@@
where $f$ is a section of $\mathcal{O}_{X}$, $D_{i}$ is a section of $\mathcal{L}$ and $\bar{D}_{i}$ is its associated image in $\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$ $(i=1,2,\dots,k)$. Sheafification of this map gives the generalized PBW map between the associated sheaves (without considering the graded quotient) [29]. In addition, if $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module then we get following results.
@@
The $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$ is a free $\mathcal{O}_{X}(U_{x})$-module for each special open set $U_{x}$. As in[19], we consider the isomorphism of cochain complexes associated with each special open set $U_{x}$ given as $Cob^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\overset{\cong}{\rightarrow}Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})).$ The isomorphism induces an isomorphism of graded vector spaces
@@
These morphisms defines cochain equivalence (or quasi isomorphism)
$Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}))\overset{\sim}{\rightarrow}(\wedge_{\mathcal{O}_{X}(U_{x})}^{\bullet}\mathcal{L}(U_{x}),0)$ on each special open set $U_{x}$ in $X$ and using that we get the isomorphism of graded vector spaces
@@
Note: We consider the presheaf of cobar resolutions as described in Section 5.1. Consider the cobar complex for $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, denote it by $\mathscr{C}ob^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))$ and the associated hypercohomology is isomorphic to the Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, i.e. $\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))$. Similar notions are applicable for the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ which we describe in the next section 5.3.
@@
Here we apply the notion of hypercohomology (Note 3.2) for the cochain complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. For each sheaf of ${\mathcal{O}_{X}}$-modules (or $\mathcal{O}_{X}$-module) $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$, we have an injective resolution (given by the flabby Godement resolution) $\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$, provides a quasi-isomorphism
@@
Notations: We denote cotensor product with $\mathcal{O}_{X}(U)$ in the category of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules as $\mathcal{O}_{X}(U)$ $\Box_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}$- and the associated cohomology groups is given by the Cotor groups (functor) as $Cotor^{\bullet}_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}(\mathcal{O}_{X}(U),-)$, for an open set $U$ of $X$. Instead of $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ we use $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ when it requires.
@@
compatible with restrictions, where $\tilde{K}^{\bullet}_{U}\mathcal{L}:=S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ and the differential $\partial_{U}:\tilde{K}^{\bullet}_{U}\mathcal{L}\rightarrow\tilde{K}^{\bullet+1}_{U}\mathcal{L}$ is described in [19], for any open set $U$ in $X$. Indeed, the dual Koszul-Rinehart complex for $\mathcal{L}(U)$ is given as $\tilde{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U):=(\tilde{K}^{\bullet}_{U}\mathcal{L},$ $\partial_{U})$.
Now, an open set $U$ is expressed by taking a partition as $\cup_{x\in X}(U\cap U_{x})$ and thus we have cochain homomorphisms of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules

### v8

@@ 3.3. Chern classes:
Lie algebroid (or Lie-Rinehart algebra) connections, curvatures and the associated Characteristic classes (e.g. Picard group, Chern class) has been studied in the context of differential geometry [10] and algebric geometry [16, 23, 24, 25]. We review and extend some of the facts in our context as follows.
@@
where $Ext^{1}(L,E,\nabla)$ is the set of equivalence classes of extensions of $L$ by the flat connection $(E,\nabla)$ [16, 25]. This result can be viewed in the algebraic geometry set up (over the affine scheme $SpecR$). We derive an analogous result for a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$ as follows.
We denote the set of equivalence classes of abelian extensions of a Lie algebroid $\mathcal{L}$ by a flat $\mathcal{L}$-connection $(\mathcal{E},\nabla)$ as $Ext^{1}(\mathcal{L},\mathcal{E},\nabla)$. An element $[\mathcal{L}^{\prime}]\in Ext^{1}(\mathcal{L},\mathcal{E},\nabla)$ represent a s.e.s. of Lie algebroids
$\mathcal{E}\rightarrow\mathcal{L}^{\prime}\rightarrow\mathcal{L}$
upto isomorphism in the usual sense, where $\mathcal{E}$ is the trivial Lie algebroid. Thus, as a local description, on the level of stalks, for each point $x\in X$, we have the s.e.s. of $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebras
$\mathcal{E}_{x}\rightarrow{\mathcal{L}^{\prime}}_{x}\rightarrow\mathcal{L}_{x}.$
Theorem 3.12.
Let $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ be a locally free Lie algebroid over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$. Then, we get a canonical isomorphism of $\mathbb{K}$-vector spaces
@@
Proof.
Given a Lie algebroid $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ over $(X,\mathcal{O}_{X})$ and an element ((((an isorphism class)))) $[\omega]\in\mathbb{H}^{2}(\mathcal{L},\mathcal{E})$ where $\omega\in\Omega^{2}_{\mathcal{L}}(\mathcal{E})$, we get a Lie algebroid extension of $\mathcal{L}$ by $\mathcal{E}$ (upto isomorphism) as follows.
Consider the $\mathcal{O}_{X}$-module $\mathcal{E}\oplus\mathcal{L}$ with the $\mathbb{K}_{X}$-Lie algebra structure given by
$[(s,D),(s^{\prime},D^{\prime})]_{\omega}:=(\nabla_{D}(s^{\prime})-\nabla_{D^{\prime}}(s)+\omega(D,D^{\prime}),[D,D^{\prime}]),$
for $s,s^{\prime}\in\mathcal{E}$ and $D,D^{\prime}\in\mathcal{L}$, together with the map
$\mathfrak{a}_{\omega}:\mathcal{L}_{\omega}\rightarrow\mathcal{T}_{X},~{}(s,D)\mapsto\mathfrak{a}(D).$
Thus, $\mathcal{L}_{\omega}:=(\mathcal{E}\oplus\mathcal{L},[\cdot,\cdot]_{\omega},\mathfrak{a}_{\omega})$ forms a Lie algebroid over $(X,\mathcal{O}_{X})$. Let $[\omega]=[\omega^{\prime}]$ for some $\omega^{\prime}\in\Omega^{2}_{\mathcal{L}}(\mathcal{E})$, i.e. $\omega-\omega^{\prime}=d\theta$ for some $\theta\in\Omega^{1}_{\mathcal{L}}(\mathcal{E})$. Then the map $\tilde{\theta}:\mathcal{L}_{\omega}\rightarrow\mathcal{L}_{\omega^{\prime}}$ defined by $\tilde{\theta}(s+D)=(s+\theta(D))+D$ forms an isomorphism of Lie algebroids. Thus, we get an abelian Lie algebroid extension (as an equivalence class) of $\mathcal{L}$ by the module $(\mathcal{E},\nabla)$ i.e. a s.e.s. of Lie algebroids where $\mathcal{E}$ is the trivial Lie algebroid
$\mathcal{E}\hookrightarrow\mathcal{L}_{\omega}\twoheadrightarrow\mathcal{L},$
and the resultant isomorphism class $[\mathcal{L}_{\omega}]\in Ext^{1}(\mathcal{L},\mathcal{E},\nabla)$.
Consider an abelian Lie algebroid extension $\mathcal{L}^{\prime}$ of $\mathcal{L}$ by $\mathcal{E}$ upto isomorphism, i.e. $[\mathcal{L}^{\prime}]\in Ext^{1}(\mathcal{L},\mathcal{E},\nabla)$ as
(8) $\mathcal{E}\rightarrow\mathcal{L}^{\prime}\rightarrow\mathcal{L},$
where $(\mathcal{E},\nabla)$ is a $\mathcal{L}$-module. Since $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module, the s.e.s. (8) splits as an $\mathcal{O}_{X}$-module, i.e. there is an isomorphism ϕ~::~italic-$\tilde{\phi}:\mathcal{L}^{\prime}\rightarrow\mathcal{E}\oplus\mathcal{L}$ for the split ((((or section map)))) ϕ::italic-$\phi:\mathcal{L}\rightarrow\mathcal{L}^{\prime}$. Thus, the possible choice for a $\mathbb{K}_{X}$-Lie bracket on $\mathcal{E}\oplus\mathcal{L}$ is given as follows:
$[(s,D),(s^{\prime},D^{\prime})]^{\prime}:=[s,s^{\prime}]^{\prime}+[D,s^{\prime}]^{\prime}-[D^{\prime},s]^{\prime}+[D,D^{\prime}]^{\prime},$
where $[s,s^{\prime}]^{\prime}=0,~{}[D,s^{\prime}]^{\prime}=\nabla_{D}(s^{\prime}),~{}[D^{\prime},s]^{\prime}=\nabla_{D^{\prime}}(s)$,  $[D,D^{\prime}]^{\prime}=\omega(D,D^{\prime})+[D,D^{\prime}]$ for some $\omega\in\Omega^{2}_{\mathcal{L}}(\mathcal{E})$ with $d(\omega)=0$,  for $s,s^{\prime}\in\mathcal{E},~{}D,D^{\prime}\in\mathcal{L}$. Let ϕ::superscriptitalic-$\phi^{\prime}:\mathcal{L}\rightarrow\mathcal{L}^{\prime}$ be another such splitting map and $\omega^{\prime}\in\Omega^{2}_{\mathcal{L}}(\mathcal{E})$ associated to the isomorphism ϕ~::~superscriptitalic-$\tilde{\phi^{\prime}}:\mathcal{L}^{\prime}\rightarrow\mathcal{E}\oplus\mathcal{L}$. Thus, using the automorphism $\tilde{\phi}^{\prime}\circ\tilde{\phi}^{-1}$ on $\mathcal{E}\oplus\mathcal{L}$ we get $\omega-\omega^{\prime}=d(\phi-{\phi}^{\prime})$. Thus, we get the required element $[\omega]\in\mathbb{H}^{2}(\mathcal{L},\mathcal{E})$ associated to the equivalence class $[\mathcal{L}^{\prime}]$.
This one-to-one correspondence between the $\mathbb{K}$-vector spaces $\mathbb{H}^{2}(\mathcal{L},\mathcal{E})$ and $Ext^{1}(\mathcal{L},\mathcal{E},\nabla)$ (with the standard vector space structure) can be extended to an isomorphism. ∎
Remark 3.13.
In particular, for $\mathcal{E}=\mathcal{O}_{X}$ with the standard $\mathcal{L}$-module structure given by the anchor map, this kind of result is proved in the holomorphic context using Čech cohomology [37].
@@
for any two sections $D,D^{\prime}$ of $\mathcal{T}_{X}$. More generally, for a $\mathcal{L}$-connection $\nabla$ on $\mathcal{E}$, one verifies that the curvature $R_{\nabla}\in\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{2}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E}))$, where $\mathcal{L}$ is a Lie algebroid.
@@
$c_{1}(\mathcal{E}):=Trace(R_{\nabla})\in\mathbb{H}^{2}(X,\mathcal{O}_{X})$
where $Trace:\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E})\rightarrow\mathcal{O}_{X}$ is the trace map.
The restriction of a $\mathcal{T}_{X}$-connection to the Lie algebroid $\mathcal{L}:=\mathcal{T}_{X}(-logY)$ ((((i.e. $\Omega_{\mathcal{L}}^{1}:=\Omega_{X}^{1}(logY)$) induces a logarithmic connection which is a meromorphic connection with poles along the divisor $Y$. Moreover, if $Y$ is a free divisor, then for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank with a $\mathcal{L}$-connection we get the first Chern class $c_{1}(\mathcal{E})\in\mathbb{H}^{2}(\mathcal{L},\mathcal{O}_{X})$. Note that a $\mathcal{L}$-connection might not exists in general. We can extend the notion of Chern classes for $\mathcal{T}_{X}$ or $\mathcal{T}_{X}(-logY)$ to arbitary Lie algebroids.
@@ 4. Logarithmic de Rham cohomology
We compute algebraic (analytic) de Rham cohomology groups of a family of non-singular varieties $Y_{t}$ for $t\in\mathbb{C}\setminus\{0\}$ and their associated singular variety $Y_{0}$ (which appears as principal divisors [29]). Thus, we derive cohomology of the Lie algebroid $\mathcal{T}_{Y_{t}}$ for $t\in\mathbb{C}\setminus\{0\}$ and study the singular case analogously. Then, we compute hypercohomology of the logarithmic de Rham complexes [27], which simplifies to the Lie algebroid cohomology for $\mathcal{T}_{X}(-logY_{t})$ [32], as we work over some free divisors $Y_{t}\subset X$ for $t\in\mathbb{C}$ [4, 32].
@@
(using the standard Lemma of Atiyah-Hodge [36]), similar results appear in [37, 4].
@@ 5. Hochschild Cohomology for a Lie algebroid over $a$-spaces
In this section, we define Hochschild cohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra (see Section 2.3), using (Hopf-)Hochschild cohomology of a left bialgebroid [19] as local descriptions. In particular, we describe the cases associated with universal enveloping algebroid and jet algebroid of a locally free Lie algebroid over an $a$-space. The cohomology groups can be computed using suitable standard complexes. We prove a version of Hochschild-Kostant-Rosenberg (HKR) theorem and dual HKR theorem. It is done by following the local counterpart as can be found in [19], where the authors proved (dual) HKR theorem for (finitely generated) projective Lie-Rinehart algebras. In particular, we present the HKR theorem for the sheaf of logarithmic differential operators and the sheaf of noncommuative differential operators.
@@ 5.1. Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra
where $C^{\bullet}(\mathcal{A}(U))=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}},b)$ is the Hochschild cochain complex of the $\mathcal{O}_{X}(U)/{\mathbb{K}}$-bialgebra $\mathcal{A}(U)$ with the differential $b:(\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}}\rightarrow(\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}}$ defined for any $a_{1},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
$b(a_{1}\otimes\cdots\otimes a_{n})=1\otimes a_{1}\otimes\cdots\otimes a_{n}+\sum^{n}_{i=1}(-1)^{i}a_{1}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}+(-1)^{n+1}a_{1}\otimes\cdots\otimes a_{n}\otimes 1~{}.$
@@
where $Cob^{\bullet}({\mathcal{A}(U)}):=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}},b^{\prime})$ is the cobar complex for $\mathcal{O}_{X}(U)$ in the category of left $\mathcal{A}(U)$-comodules with the differential $b^{\prime}:(\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}}\rightarrow(\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+2}}$ defined for any $a_{0},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
@@
Note 5.1.
The standard notion of the Hochschild chain complex is defined for an associative $\mathbb{K}$-algebra using its multiplication, and its dual complex is then considered as the Hochschild cochain complex. In our case, the algebra structure over $\mathcal{O}_{X}$ is important to consider. For an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra, we do not have an associative algebra structure over $\mathcal{O}_{X}$, but rather a co-associative algebra structure. Therefore, for an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra the co-associative algebra structure over $\mathcal{O}_{X}$ defines Hochschild cohomology using co-multiplication.
Now, we describe Hochschild (hyper)cohomology for the $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ of a (locally free) Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ (see Section 2 for these notions).
@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
The sheaf of $\mathcal{L}$-poly-vector fields on $X$ is defined as $\mathcal{T}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\wedge^{i}\mathcal{L}$. This generalizes the space of multisections of the tangent bundle (i.e. multi-vector fields) descried for different $a$-spaces occurs in geometries. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. For the case of a smooth manifold $X$, Kontsevich introduced the (sheaf of) poly-differential operators on $X$. It induces a subcomplex of the Hochschild cochain complex for $\mathcal{O}_{X}$. Its analogue in the context of Lie algebroid is the sheaf of $\mathcal{L}$-poly differential operators $\mathcal{D}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes_{\mathcal{O}_{X}}^{i}}$. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. In particular, when $\mathcal{L}=\mathcal{T}_{X}$ we denote these sheaves by the standard notions $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ respectively. The so-called HKR map is a quasi-isomorphism between the complexes associated with $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ (of differential graded $\mathbb{K}_{X}$-vector spaces). See [7] for more generalities.
@@
We know that the sheaf of symmetric algebras $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ can be viewed as universal enveloping algebroid of the Lie algebroid $\mathcal{L}$ if the $\mathcal{O}_{X}$-module $\mathcal{L}$ is equipped with zero bracket and zero anchor (on each space of sections) [29]. Recall that we have PBW sheaf homomorphism of $\mathcal{O}_{X}$-algebras given as
$\theta:\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$
@@
where $f$ is a section of $\mathcal{O}_{X}$, $D_{i}$ is a section of $\mathcal{L}$ and $\bar{D}_{i}$ is its associated image in $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ $(i=1,2,\dots,k)$. Sheafification of this map gives the generalized PBW map between the associated sheaves (without considering the graded quotient) [29]. In addition, if $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module then we get following results.
@@
The $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$ is a free $\mathcal{O}_{X}(U_{x})$-module for each special open set $U_{x}$ (3.5). As in [19], we consider the isomorphism of cochain complexes associated with each special open set $U_{x}$ given as $Cob^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\overset{\cong}{\rightarrow}Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})).$ The isomorphism induces an isomorphism of graded vector spaces
@@
These morphisms defines cochain equivalence (or quasi isomorphism) $Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}))\overset{\sim}{\rightarrow}(\wedge_{\mathcal{O}_{X}(U_{x})}^{\bullet}\mathcal{L}(U_{x}),0)$ on each special open set $U_{x}$ in $X$ and using that we get the isomorphism of graded vector spaces
@@
Here we apply the notion of hypercohomology (Note 3.1, Note 3.2) for the cochain complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. For each sheaf of ${\mathcal{O}_{X}}$-modules (or $\mathcal{O}_{X}$-module) $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$, we have an injective resolution (given by the flabby Godement resolution) $\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$, provides a quasi-isomorphism
@@
Notations: For an open set $U$ of $X$, the cotensor product with $\mathcal{O}_{X}(U)$ in the category of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules is $\mathcal{O}_{X}(U)$ $\Box_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}$- and the associated cohomology groups is given by the Cotor groups (functor) as $Cotor^{\bullet}_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}(\mathcal{O}_{X}(U),-)$. Instead of $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$, we use $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ when it requires.
@@
compatible with restrictions, where $\tilde{K}^{\bullet}_{U_{x}}\mathcal{L}:=S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})\otimes_{\mathcal{O}_{X}(U_{x})}\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$ is with the canonical $S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$-comodule structure and the differential $\partial_{U_{x}}:\tilde{K}^{\bullet}_{U_{x}}\mathcal{L}\rightarrow\tilde{K}^{\bullet+1}_{U_{x}}\mathcal{L}$ is given by the dual Koszul resolution of $\mathcal{O}(U_{x})$ as $S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$-comodules. Indeed, the dual Koszul-Rinehart complex for $\mathcal{L}(U)$ is given as $\tilde{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U):=(\tilde{K}^{\bullet}_{U}\mathcal{L},$ $\partial_{U})$, for any open set $U$ in $X$.
Now, an open set $U\subset X$ is expressed by taking a partition as $\cup_{x\in X}(U\cap U_{x})$ and thus we have cochain homomorphisms of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules